4. A 2m long fixed string has nodes at both ends when it vibrates in its fundamental mode. If the distance between this nodes is 12cm. what are the number of (1) nodes (b) anti-nodes?
1
Expert's answer
2019-11-13T09:22:01-0500

Finding the number of nodes and anti -nodes


We need to find the number of nodes and anti -nodes


Answer:


Length of the string = 2 m


a)


Nodes are the point of zero displacement


The number of nodes are

"= N + 1 = \frac {2} {0.1} + 1 = 20 + 1 = 21"


b). Antinodes are the points of maximum displacement

The number of anti-nodes are

"= N = \frac {2} {0.1} = 20"

Answer:


The number of nodes are 21

The number of anti -nodes are 20.